Global estimates of denitrification in terrestrial ecosystems are of the order 200*1012 g/yr, and approximately half of that total occurs in wetland ecosystems. If restoration efforts result in an increase in wetlands area by approximately 25%, it would be expected that losses of nitrogen from ecosystems to the atmosphere through denitrification would be closer to

275*1012 g/yr

175*1012 g/yr

225*1012 g/yr

325*1012 g/yr
Answer #1

ANSWER : C ; 225 * 1012 g/yr

Global estimates of denitrification in terrestrial ecosystems , Gt =  200*1012 g/yr

therefore half of same, ie in wetland ecosystems Gw = 100*1012 g/yr

If restoration efforts result in an increase in wetlands area by approximately 25%, then Gw1 = (1.25 * 100)*1012 g/yr

= (125) *1012 g/yr

losses of nitrogen from ecosystems to the atmosphere through denitrification = Gw + Gw1 = (225) *1012 g/yr